# Primordial Black Hole Formation via Inverted Bubble Collapse

### v1

@@ Appendix A Singlet Extension Model for EWPT
$\begin{pmatrix}s\\ \phi\end{pmatrix}=\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\begin{pmatrix}H\\ h\end{pmatrix}.$ (25)
The mass eigenstates $H$ and $h$ are identified with the heavy Higgs boson and 125 GeV Higgs boson, respectively. With the transformation (25), the masses of $H$ and $h$ and the mixing angle $\theta$ are calculated as
$m_{H}^{2}=\cos^{2}\theta M_{ss}^{2}+\sin^{2}\theta M_{hh}^{2}+\sin(2\theta)M_{sh}^{2},$ (26)
$m_{h}^{2}=\sin^{2}\theta M_{ss}^{2}+\cos^{2}\theta M_{hh}^{2}-\sin(2\theta)M_{sh}^{2},$ (27)
$\tan(2\theta)=\frac{2\mu_{\Phi S}v}{2m_{S}^{2}-v^{2}(2\lambda-\lambda_{\Phi S})}.$ (28)
@@
Though these relations and the tadpole conditions of the scalar potential (17), the potential parameters $t_{S}$,$m_{\Phi}^{2}$, $\lambda$, $m_{S}^{2}$, and $\mu_{\Phi S}$ can be expressed in terms of the physical masses and mixing parameters as
$t_{S}=-\frac{v^{2}}{2}\mu_{\Phi S},$ (29)
$m_{S}^{2}=\frac{1}{2}\left(-\sin^{2}\theta m_{h}^{2}-\cos^{2}\theta m_{H}^{2}+\lambda_{\Phi S}v^{2}\right),$ (30)
$m_{\Phi}^{2}=v^{2}\lambda,$ (31)
$\lambda=\frac{1}{2v^{2}}(\cos^{2}\theta m_{h}^{2}+\sin^{2}\theta m_{H}^{2}),$ (32)
$\mu_{\Phi S}=\frac{1}{v}(m_{H}^{2}-m_{h}^{2})\cos\theta\sin\theta.$ (33)
We choose the following seven parameters as model input parameters:
$v\;,\quad m_{h}\;,\quad m_{H}\;,\quad\theta\;,\quad\mu_{S}\;,\quad\lambda_{S}\;,\quad\lambda_{\Phi S},$ (34)
where the EW VEV and the SM-like Higgs boson mass are fixed to be $v=246.2\,{\rm GeV},m_{h}=125.09\,{\rm GeV}$.
From the potential (17), one can calculate the effective potential, which consists of the three contributions, i.e., $V_{\rm eff}(\phi,s,T)=V_{0}(\phi,s)+V_{\rm CW}(\phi,s)+V_{T}(\phi,s,T)$, where $V_{0}(\phi,s)$ is the tree-level potential, $V_{\rm CW}(\phi,s)$ is Coleman-Weinberg potential up to the one-loop level Coleman and Weinberg (1973), and $V_{T}(\phi,s,T)$ is the finite-temperature effect of the effective potential Dolan and Jackiw (1974). The tree-level contribution $V_{0}$ can be straightforwardly obtained from the potential (17) with the replacement of $(\Phi,S)\to(\phi,s)$. The remaining parts $V_{\rm CW}$ and $V_{T}$ depend on the field-dependent masses $m^{2}_{i}(\phi,s)$ of the model, for which the concrete expressions are given in, e.g., Refs. Ashoorioon and Konstandin (2009); Hashino et al. (2017); Ellis et al. (2023). Renormalization of the Coleman-Weinberg potential is performed in the $\overline{\rm MS}$ scheme, by which the renormalization scale $\mu_{R}$ is introduced in the effective potential. [^4] The Coleman-Weinberg potential can also be evaluated by applying another renormalization scheme, such as on shell-like schemes. The renormalization scheme difference of the effective potential is discussed in Refs. Chiang et al. (2019); Athron et al. (2023). In our analysis, we set $\mu_{R}=m_{t}$. We implement the thermal effects of the field-dependent masses by the replacement $m^{2}_{i}(\phi,s)\to M^{2}_{i}(\phi,s,T)=m^{2}_{i}(\phi,s)+\Pi_{i}(T)$ in the effective potential $V_{\rm eff}$ Parwani (1992), where the temperature-dependent self-energy $\Pi_{i}(T)$ is presented in, e.g., Refs. Ashoorioon and Konstandin (2009); Hashino et al. (2017); Ellis et al. (2023).
[…]
Fig. 3 shows the highly monochromatic PBH mass spectrum predicted in the singlet-extension model. It should be noted that due to the extremely narrow peak width, the PBH fraction in dark matter obtained by integrating the distribution (green dots) is about two orders of magnitude smaller than the peak height of the mass distribution. This spectrum is obtained by substituting the temperature-dependent bubble nucleation rate and $\Delta V$ calculated using CosmoTransition into Eq. (16). We have assumed that the bubbles start to shrink one Hubble time after the EW vacuum becomes the true vacuum and used $T_{c}=T_{t}/\sqrt{2}$. Here, the spectra from left to right correspond to different values of the singlet triple coupling $\mu_{S}=-246.025,-246.1,-246.175,-246.25,-246.325$ GeV, while keeping other parameters fixed as in the main text.
Figure 3: The highly monochromatic PBH mass spectrum in the singlet-extension model. The spectra from left to right are calculated with different values of the singlet triple coupling $\mu_{S}=-246.025,-246.1,-246.175,-246.25,-246.325$ GeV, while other parameters are fixed as in the main text. Note that the PBH fraction shown by the green dots is about two orders of magnitude smaller than the peak height due to the extremely narrow width.

### v3

@@ Appendix A Diagonalization of the mass matrix for the CP-even Higgs boson
$\begin{pmatrix}\phi\\ s_{1}\\ s_{2}\end{pmatrix}=R_{\theta}\begin{pmatrix}h\\ H_{1}\\ H_{2}\end{pmatrix},$ (35)
$R_{\theta}=\begin{pmatrix}c_{1}c_{2}&-s_{1}c_{2}&-s_{2}\\ s_{1}c_{3}-c_{1}s_{2}s_{3}&c_{1}c_{3}+s_{1}s_{2}s_{3}&-c_{2}s_{3}\\ c_{1}s_{2}c_{3}+s_{1}s_{3}&c_{1}s_{3}-s_{1}s_{2}c_{3}&c_{2}c_{3}\end{pmatrix},$ (36)
with the shorthanded notations $s_{i}=\sin\theta_{i}$, and $c_{i}=\cos\theta_{i}$ $(i=1,2,3)$. Thus, the masses of $H_{i}$ and $h$ are obtained by
${\rm diag}(m_{h}^{2},m_{H_{1}}^{2},m_{H_{2}}^{2})=R_{\theta}M^{2}R_{\theta}^{T}.$ (37)
@@
Through the relations (37) and the tadpole conditions of the scalar potential (16), the potential parameters $m_{\Phi}^{2}$, $m_{i}^{2}\ (i=1,2)$, $t_{1}$, $\lambda$, $\lambda_{2}$, $\lambda_{\Phi 2}$, $\mu_{\Phi 1}$ and $\mu_{12}$ can be expressed in terms of the physical masses and mixing parameters as
$m_{\Phi}^{2}=$ $\lambda v^{2}+\lambda_{\Phi 2}v_{2}^{2},$ (38)
$m_{1}^{2}=$ $-\frac{1}{2}m_{h}^{2}s^{2}_{1}c^{2}_{2}-\frac{1}{2}m_{H_{1}}^{2}(s_{1}s_{2}s_{3}+c_{1}c_{3})^{2}-\frac{1}{2}m_{H_{2}}^{2}(s_{1}s_{2}c_{3}-c_{1}s_{3})^{2}+\frac{\lambda_{\Phi 1}v^{2}}{2}+\lambda_{12}v_{2}^{2},$ (39)
$m_{2}^{2}=$ $\frac{\lambda_{\Phi 2}v^{2}}{2}+2\lambda_{2}v_{2}^{2},$ (40)
$t_{1}=$ $\frac{1}{2}\left(-\mu_{\Phi 1}v^{2}-2\mu_{12}v_{2}^{2}\right),$ (41)
$\lambda=$ $\frac{m_{h}^{2}c^{2}_{1}c^{2}_{2}}{2v^{2}}+\frac{m_{H_{1}}^{2}(s_{1}c_{3}-c_{1}s_{2}s_{3})^{2}}{2v^{2}}+\frac{m_{H_{2}}^{2}(c_{1}s_{2}c_{3}+s_{1}s_{3})^{2}}{2v^{2}},$ (42)
$\lambda_{2}=$ $\frac{m_{h}^{2}s^{2}_{2}}{8v_{2}^{2}}+\frac{m_{H_{1}}^{2}c^{2}_{2}s^{2}_{3}}{8v_{2}^{2}}+\frac{m_{H_{2}}^{2}c^{2}_{2}c^{2}_{3}}{8v_{2}^{2}},$ (43)
$\lambda_{\Phi 2}=$ $-\frac{m_{h}^{2}c_{1}s_{2}c_{2}}{2vv_{2}}+\frac{m_{H_{1}}^{2}c_{2}\left(c_{1}s_{2}s^{2}_{3}-s_{1}s_{3}c_{3}\right)}{2vv_{2}}+\frac{m_{H_{2}}^{2}c_{2}\left(c_{1}s_{2}c^{2}_{3}+s_{1}s_{3}c_{3}\right)}{2vv_{2}},$ (44)
$\mu_{\Phi 1}=$ $-\frac{m_{h}^{2}s_{1}c_{1}c^{2}_{2}}{v}+\frac{m_{H_{1}}^{2}}{8v}\Big{[}2s_{1}c_{1}\left\{2c^{2}_{2}-\left(-s^{2}_{2}+c^{2}_{2}-3\right)\left(c^{2}_{3}-s^{2}_{3}\right)\right\}-8s_{2}s_{3}c_{3}\left(c^{2}_{1}-s^{2}_{1}\right)\Big{]}$
$+\frac{m_{H_{2}}^{2}}{8v}\Big{[}8s_{2}s_{3}c_{3}\left(c^{2}_{1}-s^{2}_{1}\right)+2s_{1}c_{1}\left\{\left(-s^{2}_{2}+c^{2}_{2}-3\right)\left(c^{2}_{3}-s^{2}_{3}\right)+2c^{2}_{2}\right\}\Big{]},$ (45)
$\mu_{12}=$ $\frac{m_{h}^{2}s_{1}s_{2}c_{2}}{2v_{2}}-\frac{m_{H_{1}}^{2}c_{2}\left(s_{1}s_{2}s^{2}_{3}+c_{1}s_{3}c_{3}\right)}{2v_{2}}+\frac{m_{H_{2}}^{2}c_{2}\left(c_{1}s_{3}c_{3}-s_{1}s_{2}c^{2}_{3}\right)}{2v_{2}}\;.$ (46)
Appendix B Field-dependent masses
In the following, we list field-dependent masses of the scalar bosons in the model with two real singlet scalars. For the CP-even Higgs bosons, from the tree-level effective potential $V_{0}(\phi,s_{i})$, one obtains the field-dependent mass matrix
¯$\bar{{\cal M}}^{2}$ =(¯$=\left(\begin{array}[]{ccc}\bar{{\cal M}}_{\phi\phi}^{2}&\bar{{\cal M}}_{\phi 1}^{2}&\bar{{\cal M}}_{\phi 2}^{2}\\ \bar{{\cal M}}_{\phi 1}^{2}&\bar{{\cal M}}_{11}^{2}&\bar{{\cal M}}_{12}^{2}\\ \bar{{\cal M}}_{\phi 2}^{2}&\bar{{\cal M}}_{12}^{2}&\bar{{\cal M}}_{22}^{2}\\ \end{array}\right),$ (50)
¯$\bar{{\cal M}}^{2}_{\phi\phi}$ $=3\lambda\phi^{2}-m_{\Phi}^{2}+\lambda_{\Phi 1}s_{1}^{2}+\mu_{\Phi 1}s_{1}+\lambda_{\Phi 2}s_{2}^{2},$ (51)
¯$\bar{{\cal M}}^{2}_{\phi 1}$ $=\phi(\mu_{\Phi 1}+2\lambda_{\Phi 1}s_{1}),\ \ \ \bar{{\cal M}}^{2}_{\phi 2}=2\lambda_{\Phi 2}s_{2}\phi,$ (52)
¯$\bar{{\cal M}}_{11}^{2}$ $=\lambda_{\Phi 1}\phi^{2}-2m_{S}^{2}+12\lambda_{1}s_{1}^{2}+6\mu_{S}s_{1}+2\lambda_{12}s_{2}^{2},$ (53)
¯$\bar{{\cal M}}_{12}^{2}$ $=2s_{2}(\mu_{12}+2\lambda_{12}s_{1}),$ (54)
¯$\bar{{\cal M}}_{22}^{2}$ $=\lambda_{\Phi 2}\phi^{2}-2m_{2}^{2}+2\lambda_{12}s_{1}^{2}+2\mu_{12}s_{1}+12\lambda_{2}s_{2}^{2}.$ (55)
Diagonalizing it, one obtains field-dependent masses of the CP-even Higgs bosons, i.e., $\bar{{m}}_{h}^{2},\bar{{m}}_{H_{1}}^{2},\bar{{m}}_{H_{2}}^{2}$. The field-dependent masses of the NG bosons are given by
$\bar{m}_{G^{0}}^{2}$ $=\lambda\phi^{2}-m_{\Phi}^{2}+\lambda_{\Phi 1}s_{1}^{2}+\mu_{\Phi 1}s_{1}+\lambda_{\Phi 2}s_{2}^{2},$ (56)
$\bar{m}_{G^{\pm}}^{2}$ $=\bar{m}_{G^{0}}^{2}.$ (57)
Those of weak gauge bosons and fermions are the same as the SM [87].